OpenStudy (anonymous):

prove that the inverse tangent of x ... tan^-1(x) = 1/2 [ln ( (1+x)/(1-x) )... we can assume the 0

OpenStudy (anonymous):

is that = sign or less than sign in the equation

OpenStudy (anonymous):

tan^-1(x) equals 1/2 [ln ( (1+x)/(1-x) )

OpenStudy (anonymous):

we cud use the mean value theorem

OpenStudy (jamesj):

This relation is not true. arctan(1/sqrt(3)) of course = pi/6 but x = 1/sqrt(3) in the ln expression is not equal to this.

OpenStudy (anonymous):

cud you pls recheck your question

OpenStudy (anonymous):

omg...i'm so sorry...that wasn't tan^-1x... it should have been tanh^_1x... tanh^-1(x) = 1/2 [ln ( (1+x)/(1-x) ) hyperbolic inverse tan of x...sorry everyone!

OpenStudy (anonymous):

it s ok

OpenStudy (jamesj):

So start with x = tanh(y) using the definition of tanh and then manipulate it to solve for y

OpenStudy (anonymous):

so first wats \[\tanh x=\frac{\sinh x}{\cosh x}=\frac{ e^x-e^{-x} }{ e^x+e^{-x} }\]

OpenStudy (anonymous):

as @JamesJ suggested

OpenStudy (anonymous):

if i have ln(something/something), how do you get rid of the ln ?

OpenStudy (anonymous):

well that is the same as\[\ln \frac{a}{b}=\ln a-\ln b\] getting rid of ln completly NOT SURE

OpenStudy (jamesj):

take the exp of both sides

OpenStudy (jamesj):

to get you started ... if x = tanh(y), then x = (e^y - e^-y)/(e^y + e^-y) (e^y + e^-y)x = e^y - e^-y e^y(x-1) = - e^-y . (x+1) thus e^2y = ....

OpenStudy (anonymous):

no...its inverse tanhx

OpenStudy (jamesj):

Yes, I'm starting with tanh(y) and solving for y, because if x = tanh(y), then y = invtanh(x)

OpenStudy (anonymous):

tanh^1(x) = 1/1-x^2

OpenStudy (anonymous):

it looks like you are deriving why?

OpenStudy (anonymous):

ok...i'm trying to follow...thanks for your help btw...i'm having a hard time with this one...

OpenStudy (anonymous):

you know what Jonask, you're correct...i didn't mean to take the derivative...i'm sorry

OpenStudy (jamesj):

if x = tanh(y) then y = invtanh(x) Now if x = tanh(y), then x = (e^y - e^-y)/(e^y + e^-y) and by the algebra abovee we have e^2y = (1+x)/(1-x) thus 2y = .... and hence y = ....

OpenStudy (anonymous):

Hi James...is this leading me to prove that... tanh^-1(x) = 1/2 [ln ( (1+x)/(1-x) ) ??? I guess I don't understand why I am solving for "y"

OpenStudy (jamesj):

y is just a dummy variable. On the one hand we have y = invtanh(x) and on the other we are developing an expression for y in terms of x. If you follow what I have written for you above you will end up showing exactly that \[ y = \frac{1}{2} \ln \frac{1+x}{1-x} \]

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

@JamesJ since the relatioship between x and y is\[x^2-y^2=1\] \[\tanh y=\tan (x^2-1)\]

OpenStudy (anonymous):

is this correct to say

OpenStudy (jamesj):

no, that's not correct

OpenStudy (anonymous):

\[\tanh \sqrt{x^2-1}\]

OpenStudy (jamesj):

...at least as we have set up x and y to date. There are a few ways to prove this identity, but the way I have laid out here is the most direct.

OpenStudy (anonymous):

I need walked through this one step by step...I can't follow the logic explained above unfortunately... if I start on the left side, tanh^-1x = (e^x - e^-x)/(e^x + e^-x)... from here i need simple direction...thanks

OpenStudy (jamesj):

Ok. Let's start again. Suppose y = tanh^-1 (x) Then x = tanh(y), which means by the definition of hyperbolic tangent, tanh, that \[ x = \frac{e^y - e^{-y}}{e^y + e^{-y}} \] What you want to do now is manipulate that expression to solve for y. Does this much make sense?

OpenStudy (anonymous):

i would like to get away from the logic of soling for "y" if possible...i just wanna make the left side look like the right going from (e^x - e^-x)/(e^x + e^-x)... ...do i need to rid of the denominator somehow or multiply by a fancy form of "1" or what?

OpenStudy (anonymous):

I can't decide on my next move

OpenStudy (jamesj):

You can also do this. tanh^-1(x) = (1/2).ln( (1+x)/(1-x) ), call this second expression just s for simplicity. Then because tanh(tanh^-1 x) = x, by definition of an inverse function, it must also be the case that tanh(s) = x. So proving that tanh(s) = x is another strategy for solving the problem.

OpenStudy (jamesj):

[btw, I don't know why you didn't like the other strategy. The basic idea is very simple, isn't it? We have that y = tanh^-1(x) and we also show that y = (1/2)ln((1+x)/(1-x)) hence it must be that tanh^-1(x) = (1/2)ln((1+x)/(1-x)) But whatever floats your boat.]

OpenStudy (anonymous):

haha...i'm sorry but I just don't understand "Mathese"... I appreciate the help, I really do, but I just can't follow your logic unfortunately James...sorry but thanks

OpenStudy (anonymous):

if someone could just tell me the very next step without all the precursors and reasoning, I will probably remember what to do

OpenStudy (jamesj):

But that's the thing. This requires some thinking. It's not a grade 10 level problem.

OpenStudy (anonymous):

James, I am well adept at "thinking"...I have a 4.0 gpa over 105 credit hours in the STEM tract...I don't appreciate the grade 10 remark...did it ever occur to you that your way of explaining perhaps doesn't work for everyone? I'm not going to get in a tiff over this...thanks but you can go help someone else now